These are the guidelines that I have for the question:
Design a rectangular flower garden for the park.
The flower garden need a diagonal pathway that is 85 square root units long.
How much fencing will be needed to enclose the garden?
This is what i have tryed so far:
I know 85 square root is about 9.2, but I can't really figure out anything else.

Please please help me!!!!

Answer by richwmiller(17219)

Should I assume that all other dimensions are up to you?
As far as the sqrt(85), you will be squaring it to 85 not getting the square root.
It will be the hypotenuse of two right triangles.
Since a square is a rectangle let's make it easier for ourselves
2a^2=85
a^2=42.5
a=sqrt(42.5)
so we need
4*sqrt(42.5)+2*sqrt(5) units of fencing.
Each side of the garden is sqrt(42.5) and the diagonal need fencing on each side
Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.